OpenStudy (anonymous):

Blake has a total of 64 DVD's and CD's. If the number of CD's is eight more than three times the number of DVD's, how many DVD's does he have?

OpenStudy (johnweldon1993):

Well we have a system of equations here dvds + cds = 64 right? D + C = 64 and the numbers of cds (C) is 8 more than 3 times the dvds (D) C = 3D + 8 what would we do from here?

OpenStudy (anonymous):

Let the number of DVDs = x and the number of CDs = y According to the question x+y=64, Therefore x = 64-y y=3x+8 Now substitute x value y = 3(64-y)+8 y=192-3y+8 3y+y=200 4y=200 y=200/4 = 50 Number of DVDs = 64-50 = 14 Number of CDs = 50